An 879-kg (1943 lb) dragster, starting from rest, attains a speed of 25.4 m/s (56.9 mph) in 0.59 s. What is the average acceleration of the dragster during this time interval?
1
Expert's answer
2012-02-14T11:08:50-0500
Let'suse the following formulafor velocity of accelerated motion(taking into account that it started from rest V0=0):

V=atwhere a - acceleration, t-time

whence a=V/t
a=25.4/0.59=43